# A Variational Principle for Extended Irreversible Thermodynamics: Heat Conducting Viscous Fluids

### v1

@@ 3.2 Energy and entropy balance in the material description
$\widehat{\mathfrak{E}}=\frac{\partial\widehat{\mathfrak{L}}}{\partial\dot{\boldsymbol{\varphi}}}\cdot\dot{\boldsymbol{\varphi}}-\widehat{\mathfrak{L}}.$
@@
$\frac{d}{dt}\widehat{\mathfrak{E}}=\operatorname{DIV}\Big{(}\boldsymbol{P}\cdot\dot{\boldsymbol{\varphi}}-\frac{\partial\widehat{\mathfrak{L}}}{\partial\nabla\boldsymbol{\varphi}}\cdot\dot{\boldsymbol{\varphi}}\Big{)}-\boldsymbol{P}:\nabla\dot{\boldsymbol{\varphi}}-\frac{\partial\widehat{\mathfrak{L}}}{\partial S}\dot{S}-\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{S}}\dot{\boldsymbol{S}}-\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{Q}}\dot{\boldsymbol{Q}}.$
@@ 3.3 Variational formulation in the spatial description
Material covariance.
In order to obtain the spatial description, we shall extend the notion of material covariance given in (2.9) to the class of nonequilibrium Lagrangians. We say that $\widehat{\mathfrak{L}}\big{(}\boldsymbol{\varphi},\dot{\boldsymbol{\varphi}},\nabla{\boldsymbol{\varphi}},\varrho,S,\boldsymbol{S},\boldsymbol{Q}\big{)}$ is material covariant if it satisfies
@@
$(\boldsymbol{\psi}^{*}\boldsymbol{S})^{AB}=(\boldsymbol{S}^{CD}(\boldsymbol{\psi}^{-1})_{,C}^{A}(\boldsymbol{\psi}^{-1})_{,D}^{B})\circ\boldsymbol{\psi}\quad\text{and}\quad(\boldsymbol{\psi}^{*}\boldsymbol{Q})^{A}=(\boldsymbol{Q}^{B}(\boldsymbol{\psi}^{-1})_{,B}^{A})\circ\boldsymbol{\psi}.$
@@
Nonequilbrium Lagrangian in the spatial description.
From the material covariance property (2.9) we can associate to $\widehat{\mathfrak{L}}$ a nonequilibrium Lagrangian density $\widehat{\ell}$ in the spatial description defined by
@@
Standard Lagrangian densities.
For the Lagrangian (3.7), since the first term is clearly material covariant, the property (3.13) is equivalent to the following condition on the material nonequilibrium internal energy function
^$\widehat{\mathscr{E}}\big{(}\nabla(\boldsymbol{\varphi}\circ\boldsymbol{\psi}),(\varrho\circ\boldsymbol{\psi})J_{\boldsymbol{\psi}},(S\circ\boldsymbol{\psi})J_{\boldsymbol{\psi}},(\boldsymbol{\psi}^{*}\boldsymbol{S})J_{\boldsymbol{\psi}},(\boldsymbol{\psi}^{*}\boldsymbol{Q})J_{\boldsymbol{\psi}}\big{)}=\widehat{\mathscr{E}}\big{(}\nabla{\boldsymbol{\varphi}},\varrho,S,\boldsymbol{S},\boldsymbol{Q}\big{)}\circ\boldsymbol{\psi}$
@@
^$\widehat{\mathscr{E}}\big{(}\nabla{\boldsymbol{\varphi}},\varrho,S,\boldsymbol{S},\boldsymbol{Q}\big{)}\varrho=\big{(}\widehat{\varepsilon}\,(\rho,s,\boldsymbol{\sigma},\boldsymbol{q})\circ\boldsymbol{\varphi}\big{)}J_{\boldsymbol{\varphi}}$
@@
^$\widehat{\ell}(\boldsymbol{u},\rho,s,\boldsymbol{\sigma},\boldsymbol{q})=\frac{1}{2}\rho|\boldsymbol{u}|^{2}-\widehat{\varepsilon}\,(\rho,s,\boldsymbol{\sigma},\boldsymbol{q}).$ (3.17)
Derivation of the variational formulation.
We derive here the spatial version of the variational formulation (3.1)–(3.3). This is achieved by converting each of the three conditions, as done in (i),(ii),(iii) below.
@@
(ii) To get the spatial version of the phenomenological constraint (3.2), we note that when $\boldsymbol{S}(t,\mathbf{X})$ and $\boldsymbol{Q}(t,\mathbf{X})$ are related to $\boldsymbol{\sigma}(t,\mathbf{x})$ and $\boldsymbol{q}(t,\mathbf{x})$ via the formulas (3.15) for some time dependent diffeomorphism $\boldsymbol{\varphi}(t,\mathbf{X})$, we have the following identities
$\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{S}}:\dot{\boldsymbol{S}}$ =((^$=\bigg{(}\Big{(}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{t}\boldsymbol{\sigma}\Big{)}\circ\boldsymbol{\varphi}\bigg{)}J_{\boldsymbol{\varphi}}$ (3.18)
$\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{Q}}\cdot\dot{\boldsymbol{Q}}$ =((^$=\bigg{(}\Big{(}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{t}\boldsymbol{q}\Big{)}\circ\boldsymbol{\varphi}\bigg{)}J_{\boldsymbol{\varphi}},$ (3.19)
@@
with (£$(\pounds_{\boldsymbol{u}}\boldsymbol{\sigma})^{ab}=\boldsymbol{u}^{c}\partial_{c}\boldsymbol{\sigma}^{ab}-\boldsymbol{\sigma}^{cb}\partial_{c}\boldsymbol{u}^{a}-\boldsymbol{\sigma}^{ac}\partial_{c}\boldsymbol{u}^{b}$ and (£$(\pounds_{\boldsymbol{u}}\boldsymbol{q})^{a}=\boldsymbol{u}^{c}\partial_{c}\boldsymbol{q}^{a}-\boldsymbol{q}^{c}\partial_{c}\boldsymbol{u}^{a}$ the Lie derivative of a contravariant tensor field and a vector field.
To prove (3.18)–(3.19) we first take the partial derivative of the equality (3.14) with respect to $\boldsymbol{Q}$ in some direction $\boldsymbol{V}$, and get
$\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{Q}}\cdot\boldsymbol{V}=\Big{(}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\boldsymbol{\varphi}_{*}\boldsymbol{V}J_{\boldsymbol{\varphi}^{-1}}\Big{)}\circ\boldsymbol{\varphi}J_{\boldsymbol{\varphi}}=\Big{(}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\boldsymbol{\varphi}_{*}\boldsymbol{V}\Big{)}\circ\boldsymbol{\varphi}=\boldsymbol{\varphi}^{*}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\boldsymbol{V}.$
@@
$\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{S}}:\boldsymbol{W}=\Big{(}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\boldsymbol{\varphi}_{*}\boldsymbol{W}J_{\boldsymbol{\varphi}^{-1}}\Big{)}\circ\boldsymbol{\varphi}J_{\boldsymbol{\varphi}}=\Big{(}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\boldsymbol{\varphi}_{*}\boldsymbol{W}\Big{)}\circ\boldsymbol{\varphi}=\boldsymbol{\varphi}^{*}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\boldsymbol{W}.$
@@
$\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{S}}:\dot{\boldsymbol{S}}=\boldsymbol{\varphi}^{*}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}\cdot(\boldsymbol{\varphi}^{*}\mathfrak{D}_{t}\boldsymbol{\sigma})J_{\boldsymbol{\varphi}}=\bigg{(}\Big{(}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{t}\boldsymbol{\sigma}\Big{)}\circ\boldsymbol{\varphi}\bigg{)}J_{\boldsymbol{\varphi}}$
@@
$\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{Q}}\cdot\dot{\boldsymbol{Q}}=\boldsymbol{\varphi}^{*}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot(\boldsymbol{\varphi}^{*}\mathfrak{D}_{t}\boldsymbol{q})J_{\boldsymbol{\varphi}}=\bigg{(}\Big{(}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{t}\boldsymbol{q}\Big{)}\circ\boldsymbol{\varphi}\bigg{)}J_{\boldsymbol{\varphi}}$
@@
$\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{S}}:\delta\boldsymbol{S}$ =((^$=\bigg{(}\Big{(}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{\delta}\boldsymbol{\sigma}\Big{)}\circ\boldsymbol{\varphi}\bigg{)}J_{\boldsymbol{\varphi}}$ (3.24)
$\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{Q}}\cdot\delta\boldsymbol{Q}$ =((^$=\bigg{(}\Big{(}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{\delta}\boldsymbol{q}\Big{)}\circ\boldsymbol{\varphi}\bigg{)}J_{\boldsymbol{\varphi}},$ (3.25)
@@
The variational formulation.
By using (3.18), (3.19), (3.24), (3.25), we can convert the variational formulation (3.1)–(3.3) into the spatial description and get the following.
@@
^$\frac{\partial\widehat{\ell}}{\partial s}\bar{D}_{t}\varsigma+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{t}\boldsymbol{\sigma}+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{t}\boldsymbol{q}=-\boldsymbol{\sigma}:\nabla\boldsymbol{u}+\boldsymbol{j}\cdot\nabla D_{t}\gamma$ (3.29)
@@
^$\frac{\partial\widehat{\ell}}{\partial s}\bar{D}_{\delta}\varsigma+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{\delta}\boldsymbol{\sigma}+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{\delta}\boldsymbol{q}=-\boldsymbol{\sigma}:\nabla\boldsymbol{\zeta}+\boldsymbol{j}\cdot\nabla D_{\delta}\gamma,$ (3.30)
@@
Derivation of the critical conditions.
By taking the variations in (3.28), we get
@@
^$\frac{\partial\widehat{\ell}}{\partial s}=-D_{t}\gamma.$ (3.33)
@@
+^$\hskip 56.9055pt+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\delta\boldsymbol{\sigma}+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\delta\boldsymbol{q}-\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{\delta}\boldsymbol{\sigma}-\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{\delta}\boldsymbol{q}\Big{]}{\rm d}^{3}\mathbf{x}\,{\rm d}t=0,$
@@
$\boldsymbol{\tau}_{\boldsymbol{q}}:$ =(^$=\Big{(}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\boldsymbol{q}\Big{)}\delta-\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\otimes\boldsymbol{q},\qquad(\boldsymbol{\tau}_{\boldsymbol{q}})^{a}_{b}=\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}^{c}}\boldsymbol{q}^{c}\delta^{a}_{b}-\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}^{b}}\cdot\boldsymbol{q}^{a},$ (3.39)
$\boldsymbol{\tau}_{\boldsymbol{\sigma}}:$ =(^$=\Big{(}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\boldsymbol{\sigma}\Big{)}\delta-2\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}\cdot\boldsymbol{\sigma},\qquad(\boldsymbol{\tau}_{\boldsymbol{\sigma}})^{a}_{b}=\Big{(}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}^{cd}}\boldsymbol{\sigma}^{cd}\Big{)}\delta_{b}^{a}-2\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}^{bd}}\boldsymbol{\sigma}^{ad}.$
@@
$\left\{\begin{array}[]{l}\vspace{0.2cm}\displaystyle\partial_{t}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{u}}+\pounds_{\boldsymbol{u}}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{u}}=\rho\nabla\frac{\partial\widehat{\ell}}{\partial\rho}+s\nabla\frac{\partial\widehat{\ell}}{\partial s}-\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\nabla\boldsymbol{\sigma}-\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\nabla\boldsymbol{q}+\operatorname{div}(\boldsymbol{\sigma}+\boldsymbol{\tau}_{\boldsymbol{\sigma}}+\boldsymbol{\tau}_{\boldsymbol{q}})\\ \vspace{0.2cm}\displaystyle-\frac{\partial\widehat{\ell}}{\partial s}(\bar{D}_{t}s+\operatorname{div}\boldsymbol{j})=\boldsymbol{\sigma}:\nabla\boldsymbol{u}+\boldsymbol{j}\cdot\nabla\frac{\partial\widehat{\ell}}{\partial s}+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{t}\boldsymbol{\sigma}+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{t}\boldsymbol{q}\\ \displaystyle\bar{D}_{t}\rho=0.\end{array}\right.$ (3.41)
@@
$D_{t}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{u}}+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{u}}\operatorname{div}\boldsymbol{u}=\operatorname{div}\bigg{[}\Big{(}\frac{\partial\widehat{\ell}}{\partial\rho}\rho+\frac{\partial\widehat{\ell}}{\partial s}s+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\boldsymbol{q}+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\boldsymbol{\sigma}-\ell\Big{)}\delta-\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\otimes\boldsymbol{q}-2\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}\cdot\boldsymbol{\sigma}+\boldsymbol{\sigma}\bigg{]},$

### v2

@@ 3.2 Energy and entropy balance in the material description
$\widehat{\mathfrak{E}}=\frac{\partial\widehat{\mathfrak{L}}}{\partial\dot{\boldsymbol{\varphi}}}\cdot\dot{\boldsymbol{\varphi}}-\widehat{\mathfrak{L}}$. (3.11)
@@
$\frac{d}{dt}\widehat{\mathfrak{E}}=\operatorname{DIV}\Big{(}\boldsymbol{P}\cdot\dot{\boldsymbol{\varphi}}-\frac{\partial\widehat{\mathfrak{L}}}{\partial\nabla\boldsymbol{\varphi}}\cdot\dot{\boldsymbol{\varphi}}\Big{)}-\boldsymbol{P}:\nabla\dot{\boldsymbol{\varphi}}-\frac{\partial\widehat{\mathfrak{L}}}{\partial S}\dot{S}-\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{S}}\dot{\boldsymbol{S}}-\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{Q}}\dot{\boldsymbol{Q}}.$ (3.13)
@@ 3.3 Variational formulation in the spatial description
Material covariance. In order to obtain the spatial description, we shall extend the notion of material covariance given in (2.11) to the class of nonequilibrium Lagrangians. We say that $\widehat{\mathfrak{L}}\big{(}\boldsymbol{\varphi},\dot{\boldsymbol{\varphi}},\nabla{\boldsymbol{\varphi}},\varrho,S,\boldsymbol{S},\boldsymbol{Q}\big{)}$ is material covariant if it satisfies
@@
$(\boldsymbol{\psi}^{*}\boldsymbol{S})^{AB}=(\boldsymbol{S}^{CD}(\boldsymbol{\psi}^{-1})_{,C}^{A}(\boldsymbol{\psi}^{-1})_{,D}^{B})\circ\boldsymbol{\psi}\quad\text{and}\quad(\boldsymbol{\psi}^{*}\boldsymbol{Q})^{A}=(\boldsymbol{Q}^{B}(\boldsymbol{\psi}^{-1})_{,B}^{A})\circ\boldsymbol{\psi}.$ (3.19)
@@
Nonequilbrium Lagrangian in the spatial description. From the material covariance property (2.11) we can associate to $\widehat{\mathfrak{L}}$ a nonequilibrium Lagrangian density $\widehat{\ell}$ in the spatial description defined by
@@
Standard Lagrangian densities. For the Lagrangian (3.8), since the first term is clearly material covariant, the property (3.18) is equivalent to the following condition on the material nonequilibrium internal energy function
$\widehat{\mathscr{E}}\big{(}\nabla(\boldsymbol{\varphi}\circ\boldsymbol{\psi}),(\varrho\circ\boldsymbol{\psi})J_{\boldsymbol{\psi}},(S\circ\boldsymbol{\psi})J_{\boldsymbol{\psi}},(\boldsymbol{\psi}^{*}\boldsymbol{S})J_{\boldsymbol{\psi}},(\boldsymbol{\psi}^{*}\boldsymbol{Q})J_{\boldsymbol{\psi}}\big{)}=\widehat{\mathscr{E}}\big{(}\nabla{\boldsymbol{\varphi}},\varrho,S,\boldsymbol{S},\boldsymbol{Q}\big{)}\circ\boldsymbol{\psi}$ (3.23)
@@
$\widehat{\mathscr{E}}\big{(}\nabla{\boldsymbol{\varphi}},\varrho,S,\boldsymbol{S},\boldsymbol{Q}\big{)}\varrho=\big{(}\widehat{\varepsilon}\,(\rho,s,\boldsymbol{\sigma},\boldsymbol{q})\circ\boldsymbol{\varphi}\big{)}J_{\boldsymbol{\varphi}}$ (3.24)
@@
$\widehat{\ell}(\boldsymbol{u},\rho,s,\boldsymbol{\sigma},\boldsymbol{q})=\frac{1}{2}\rho|\boldsymbol{u}|^{2}-\widehat{\varepsilon}\,(\rho,s,\boldsymbol{\sigma},\boldsymbol{q}).$ (3.25)
Derivation of the variational formulation. We derive here the spatial version of the variational formulation (3.2)–(3.4). This is achieved by converting each of the three conditions, as done in (i),(ii),(iii) below.
@@
(ii) To get the spatial version of the phenomenological constraint (3.3), we note that when $\boldsymbol{S}(t,\mathbf{X})$ and $\boldsymbol{Q}(t,\mathbf{X})$ are related to $\boldsymbol{\sigma}(t,\mathbf{x})$ and $\boldsymbol{q}(t,\mathbf{x})$ via the formulas (3.21) for some time dependent diffeomorphism $\boldsymbol{\varphi}(t,\mathbf{X})$, we can show the following identities
$\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{S}}:\dot{\boldsymbol{S}}$ $=\bigg{(}\Big{(}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{t}\boldsymbol{\sigma}\Big{)}\circ\boldsymbol{\varphi}\bigg{)}J_{\boldsymbol{\varphi}}$ (3.26)
$\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{Q}}\cdot\dot{\boldsymbol{Q}}$ $=\bigg{(}\Big{(}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{t}\boldsymbol{q}\Big{)}\circ\boldsymbol{\varphi}\bigg{)}J_{\boldsymbol{\varphi}},$ (3.27)
@@
with $(\pounds_{\boldsymbol{u}}\boldsymbol{\sigma})^{ab}=\boldsymbol{u}^{c}\partial_{c}\boldsymbol{\sigma}^{ab}-\boldsymbol{\sigma}^{cb}\partial_{c}\boldsymbol{u}^{a}-\boldsymbol{\sigma}^{ac}\partial_{c}\boldsymbol{u}^{b}$ and $(\pounds_{\boldsymbol{u}}\boldsymbol{q})^{a}=\boldsymbol{u}^{c}\partial_{c}\boldsymbol{q}^{a}-\boldsymbol{q}^{c}\partial_{c}\boldsymbol{u}^{a}$ the Lie derivative of a contravariant tensor field and a vector field. To prove (3.26)–(3.27) we first take the partial derivative of the equality (3.20) with respect to $\boldsymbol{Q}$ in some direction $\boldsymbol{V}$, and get
$\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{Q}}\cdot\boldsymbol{V}=\Big{(}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\boldsymbol{\varphi}_{*}\boldsymbol{V}J_{\boldsymbol{\varphi}^{-1}}\Big{)}\circ\boldsymbol{\varphi}J_{\boldsymbol{\varphi}}=\Big{(}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\boldsymbol{\varphi}_{*}\boldsymbol{V}\Big{)}\circ\boldsymbol{\varphi}=\boldsymbol{\varphi}^{*}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\boldsymbol{V}.$ (3.30)
@@
$\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{S}}:\boldsymbol{W}=\Big{(}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\boldsymbol{\varphi}_{*}\boldsymbol{W}J_{\boldsymbol{\varphi}^{-1}}\Big{)}\circ\boldsymbol{\varphi}J_{\boldsymbol{\varphi}}=\Big{(}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\boldsymbol{\varphi}_{*}\boldsymbol{W}\Big{)}\circ\boldsymbol{\varphi}=\boldsymbol{\varphi}^{*}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\boldsymbol{W}.$ (3.31)
@@
$\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{S}}:\dot{\boldsymbol{S}}=\boldsymbol{\varphi}^{*}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}\cdot(\boldsymbol{\varphi}^{*}\mathfrak{D}_{t}\boldsymbol{\sigma})J_{\boldsymbol{\varphi}}=\bigg{(}\Big{(}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{t}\boldsymbol{\sigma}\Big{)}\circ\boldsymbol{\varphi}\bigg{)}J_{\boldsymbol{\varphi}}$ (3.34)
@@
$\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{Q}}\cdot\dot{\boldsymbol{Q}}=\boldsymbol{\varphi}^{*}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot(\boldsymbol{\varphi}^{*}\mathfrak{D}_{t}\boldsymbol{q})J_{\boldsymbol{\varphi}}=\bigg{(}\Big{(}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{t}\boldsymbol{q}\Big{)}\circ\boldsymbol{\varphi}\bigg{)}J_{\boldsymbol{\varphi}}$ (3.35)
@@
$\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{S}}:\delta\boldsymbol{S}$ $=\bigg{(}\Big{(}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{\delta}\boldsymbol{\sigma}\Big{)}\circ\boldsymbol{\varphi}\bigg{)}J_{\boldsymbol{\varphi}}$ (3.36)
$\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{Q}}\cdot\delta\boldsymbol{Q}$ $=\bigg{(}\Big{(}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{\delta}\boldsymbol{q}\Big{)}\circ\boldsymbol{\varphi}\bigg{)}J_{\boldsymbol{\varphi}},$ (3.37)
@@
The variational formulation. By using (3.26), (3.27), (3.36), (3.37), we can convert the variational formulation (3.2)–(3.4) into the spatial description and get the following.
@@
$\frac{\partial\widehat{\ell}}{\partial s}\bar{D}_{t}\varsigma+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{t}\boldsymbol{\sigma}+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{t}\boldsymbol{q}=-\boldsymbol{\sigma}:\nabla\boldsymbol{u}+\boldsymbol{j}\cdot\nabla D_{t}\gamma$ (3.41)
@@
$\frac{\partial\widehat{\ell}}{\partial s}\bar{D}_{\delta}\varsigma+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{\delta}\boldsymbol{\sigma}+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{\delta}\boldsymbol{q}=-\boldsymbol{\sigma}:\nabla\boldsymbol{\zeta}+\boldsymbol{j}\cdot\nabla D_{\delta}\gamma,$ (3.42)
@@
Derivation of the critical conditions. By taking the variations in (3.40), we get
@@
$\frac{\partial\widehat{\ell}}{\partial s}=-D_{t}\gamma.$ (3.45)
@@
$\hskip 56.9055pt+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\delta\boldsymbol{\sigma}+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\delta\boldsymbol{q}-\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{\delta}\boldsymbol{\sigma}-\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{\delta}\boldsymbol{q}\Big{]}{\rm d}^{3}\mathbf{x}\,{\rm d}t=0,$
@@
$\boldsymbol{\tau}_{\boldsymbol{q}}:$ $=\Big{(}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\boldsymbol{q}\Big{)}\delta-\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\otimes\boldsymbol{q},$ $\qquad(\boldsymbol{\tau}_{\boldsymbol{q}})^{a}_{b}$ $=\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}^{c}}\boldsymbol{q}^{c}\delta^{a}_{b}-\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}^{b}}\cdot\boldsymbol{q}^{a},$ (3.51)
$\boldsymbol{\tau}_{\boldsymbol{\sigma}}:$ $=\Big{(}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\boldsymbol{\sigma}\Big{)}\delta-2\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}\cdot\boldsymbol{\sigma},$ $\qquad(\boldsymbol{\tau}_{\boldsymbol{\sigma}})^{a}_{b}$ $=\Big{(}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}^{cd}}\boldsymbol{\sigma}^{cd}\Big{)}\delta_{b}^{a}-2\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}^{bd}}\boldsymbol{\sigma}^{ad}.$
@@
$\left\{\begin{array}[]{l}\vskip 5.69046pt\displaystyle\partial_{t}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{u}}+\pounds_{\boldsymbol{u}}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{u}}=\rho\nabla\frac{\partial\widehat{\ell}}{\partial\rho}+s\nabla\frac{\partial\widehat{\ell}}{\partial s}-\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\nabla\boldsymbol{\sigma}-\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\nabla\boldsymbol{q}+\operatorname{div}(\boldsymbol{\sigma}+\boldsymbol{\tau}_{\boldsymbol{\sigma}}+\boldsymbol{\tau}_{\boldsymbol{q}})\\ \vskip 5.69046pt\displaystyle-\frac{\partial\widehat{\ell}}{\partial s}(\bar{D}_{t}s+\operatorname{div}\boldsymbol{j})=\boldsymbol{\sigma}:\nabla\boldsymbol{u}+\boldsymbol{j}\cdot\nabla\frac{\partial\widehat{\ell}}{\partial s}+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{t}\boldsymbol{\sigma}+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{t}\boldsymbol{q}\\ \displaystyle\bar{D}_{t}\rho=0.\end{array}\right$. (3.53)
@@
$D_{t}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{u}}+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{u}}\operatorname{div}\boldsymbol{u}=\operatorname{div}\bigg{[}\Big{(}\frac{\partial\widehat{\ell}}{\partial\rho}\rho+\frac{\partial\widehat{\ell}}{\partial s}s+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\boldsymbol{q}+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\boldsymbol{\sigma}-\ell\Big{)}\delta-\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\otimes\boldsymbol{q}-2\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}\cdot\boldsymbol{\sigma}+\boldsymbol{\sigma}\bigg{]},$ (3.54)